OpenStudy (anonymous):

What is the fifth term of a sequence whose first term is 5 and whose common ratio is 3? 1,125?

OpenStudy (anonymous):

an = ar^(n-1)

OpenStudy (anonymous):

=5*3^(5-1) = 5.3^4

OpenStudy (anonymous):

\[\huge a_n = a \cdot r^{n-1}\] a = 5 r = 3 and n = 5 here...

OpenStudy (anonymous):

i think it is 405

OpenStudy (anonymous):

@waterineyes wat do u think

OpenStudy (anonymous):

Yes yahoo.com..

OpenStudy (anonymous):

httankyou!!